# Comparison Principles for Stochastic Volterra equations

### v1

@@ 2. Non-negativity preserving case
admits a unique strong continuous solution for $i=1,2$. Since completely monotone kernels are non-negativity preserving and satisfy (3), see [6, Theorem 2.3], we may apply steps 1 and 2 to obtain $\mathbb{P}[X^{1,H}_{T}\leq X^{2,H}_{T}]=1$ for each $T>0$. Next we prove that
@@
where $R$ denotes the resolvent of the second kind of $c(K^{H})^{2}$, where $c$ is some large enough constant. Finally, the first term can be bounded by
$\int_{0}^{T}|K(s)-K^{H}(s)|^{2+\eta}\,\mathrm{d}s$ $\leq\int_{0}^{T}\left(\int_{[H,\infty)}\mathrm{e}^{-\rho s}\,\mu(\mathrm{d}\rho)\right)^{\frac{\eta}{2+\eta}}\,\mathrm{d}s$
@@
Note that the case where $g_{i}(t)=x_{i}$ for $i=1,2$ was already covered in [6]. Since the fractional Riemann-Liouville kernel satisfies Assumption 2.2 and is completely monotone, Theorem 2.4 can be applied to this case. Our next example shows that our comparison principle may also hold when $g_{1},g_{2}$ are not admissible.
@@
Let $X^{1},X^{2}$ be fractional Ornstein-Uhlenbeck processes given as the unique strong solutions of (2), i.e.,
@@
Since $E_{\alpha,\alpha}(\beta t^{\alpha})\geq 0$ for any $\beta,t$, the comparison principle is satisfied if and only if $E_{\alpha,\gamma}(\beta t^{\alpha})\geq 0$ for all $t\geq 0$. If $\beta\geq 0$, then $E_{\alpha,\gamma}(\beta t^{\alpha})\geq 0$ and nothing needs to be shown. Suppose that $\beta<0$. If $\gamma\geq\alpha$, then $t\longmapsto E_{\alpha,\gamma}(-t^{\alpha})$ is completely monotone, see [24, Section 4.9.2], and hence the comparison principle holds. In this case $g_{1},g_{2}$ are admissible, see Example 1.2. Conversely, suppose that $\gamma<\alpha$. Then $(0,\infty)\ni t\longmapsto t^{\gamma-1}E_{\alpha,\gamma}(-|\beta|t^{\alpha})$ is integrable and its Laplace transform is given by $\int_{0}^{\infty}t^{\gamma-1}E_{\alpha,\gamma}(-|\beta|t^{\alpha})\mathrm{e}^{-st}\,\mathrm{d}s=\frac{s^{\alpha-\gamma}}{s^{\alpha}+|\beta|}$ for $s\geq 0$. Evaluating this in $s=0$, shows that $\int_{0}^{\infty}t^{\gamma-1}E_{\alpha,\gamma}(-|\beta|t^{\alpha})\,\mathrm{d}s=0$ and hence $(0,\infty)\ni t\longmapsto E_{\alpha,\gamma}(-|\beta|t^{\alpha})$ needs to attain negative values. Consequently, the comparison principle fails to hold. ∎
Next, let us examine cases where the coefficients $b_{1},b_{2},\sigma$ are not Lipschitz continuous. Since, in this generality, the weak existence of solutions is not guaranteed from the literature, below we complement the latter by proving the weak existence of continuous solutions when the function $g$ may be singular in $t=0$. Our proof follows the typical tightness arguments, however, in contrast to existing arguments such as [33, 4], here we introduce an auxiliary weighted space of continuous functions to control the order of singularity of the function $g$.
@@
then the stochastic Volterra equation
(10) $X_{t}=g(t)+\int_{0}^{t}K(t-s)b(s,X_{s})\,\mathrm{d}s+\int_{0}^{t}K(t-s)\sigma(s,X_{s})\,\mathrm{d}B_{s},\qquad t>0$
admits a continuous weak solution.
@@
Following [33, Lemma 3.6], we define for $f\in\{b,\sigma\}$ a sequence of functions
(11) $f_{n}(t,x)=\psi_{n}(x)\int_{\mathbb{R}}f(t,x-y)\varphi_{n}(y)\,\mathrm{d}y,\quad,n\in\mathbb{N},$
where $\psi_{n}\colon\mathbb{R}\longrightarrow[0,1]$ with $\operatorname{supp}\psi_{n}\subset[-(n+1),n+1]$ and $\psi_{n}(x)=1$ for $x\in[-n,n]$, where $\varphi(y)=\frac{1}{c_{n}}(1-y^{2})^{n}\mathbbm{1}_{[-1,1]}(y)$ with $c_{n}=\int_{[-1,1]}(1-y^{2})^{n}\,\mathrm{d}y$. Then $(f_{n})_{n\in\mathbb{N}}$ is a sequence of Lipschitz continuous functions in the second variable and $\lim_{n\longrightarrow\infty}f_{n}(t,\cdot)=f(t,\cdot)$ uniformly on compacts for all $t\geq 0$. Moreover, the following growth condition is satisfied uniformly in $n$,
(12) $|b_{n}(t,x)|+|\sigma_{n}(t,x)|\lesssim C(1+|x|)^{\xi},\quad\text{a.a. }t\geq 0,x\in\mathbb{R},$
@@
holds for some $\theta>0$, $p\geq 2$ and a constant $c$ that only depends on $C$, $T$, $K$, $p$. By applying Kolmogorov’s tightness criterion (see e.g. [34, Chapter XIII, Theorem 1.8]), we conclude that $(X^{n}-g)_{n\in\mathbb{N}}$ is a tight sequence of processes and hence has a weakly convergent subsequence on the path space $C(\mathbb{R}_{+})$ with limit denoted by ¯$\overline{\mathbb{P}}_{g}$. By abuse of notation, we denote this subsequence also by $(X^{n}-g)_{n\in\mathbb{N}}$. Let $C(\mathbb{R}_{+})_{0}\subset C(\mathbb{R}_{+})$ be the space of continuous functions that vanish at $t=0$. Since $\mathcal{L}(X^{n}-g)(C(\mathbb{R}_{+})_{0})=1$ it follows by the Portmanteau theorem that
@@
Next, to prove weak convergence of $(X^{n})_{n\in\mathbb{N}}$, let $C_{\kappa}([0,T])$ with $\kappa>0$ be the Banach space of continuous functions $f\colon(0,T]\longrightarrow\mathbb{R}$ with finite norm $\|f\|_{\kappa,T}\coloneqq\sup_{t\in(0,T]}t^{\kappa}|f(t)|$ such that $\lim_{t\downarrow 0}t^{\kappa}f(t)$ exists. Note that $\iota\colon C_{\kappa}([0,T])\longrightarrow C([0,T])$ defined by $\iota(f)(t)=t^{\kappa}f(t)$ satisfies $\|\iota(f)\|_{[0,T]}=\|f\|_{\kappa,T}$, and hence is an isometric isomorphism, whence $C_{\kappa}([0,T])$ is a polish space. Finally, we let $C_{\kappa}(\mathbb{R}_{+})\coloneqq\cap_{k\in\mathbb{N}}C_{\kappa}([0,k])$ be equipped with the metric
@@
By assumption (9), $g\in C_{\kappa}(\mathbb{R}_{+})$ for $\kappa>\delta$. Hence $\Phi\colon C(\mathbb{R}_{+})_{0}\longrightarrow C_{\kappa}(\mathbb{R}_{+})$ given by $\Phi(f)=f+g$ defines a continuous mapping. Thus, $(\Phi(X^{n}-g))_{n\in\mathbb{N}}=(X^{n})_{n\in\mathbb{N}}$ is weakly convergent on $C_{\kappa}(\mathbb{R}_{+})$ with weak limit denoted by $\overline{\mathbb{P}}$. Note that ¯$\overline{\mathbb{P}}_{g},\overline{\mathbb{P}}$ are related by
(13) ¯$\overline{\mathbb{P}}_{g}\circ\Phi^{-1}=\overline{\mathbb{P}}.$
It remains to show that $\overline{\mathbb{P}}$ determines a weak solution of (10). Let $L$ be the resolvent of the first kind of $K$111i.e. a locally finite (positive) measure $L$ such that $L\ast K=K\ast L=1$. The latter exists since $K\neq 0$ is non-negative and non-increasing, see [25, Theorem 5.5.5].. Let $(M_{t}^{n})_{t\geq 0}$ be the continuous $\mathbb{P}$-martingale for the filtration $\mathbb{F}$ given by
@@
where the continuity is guaranteed by [25, Corollary 3.6.2 (iii)]. Since $X^{n}-g\in C(\mathbb{R}_{+})_{0}$, by continuous mapping theorem combined with (13), we obtain
@@
Next, we show that $(\mathcal{M}^{\Phi}_{t})_{t\geq 0}$ is a continuous $\overline{\mathbb{P}}$-martingale for the natural filtration generated by coordinate process on $C_{\kappa}(\mathbb{R}_{+})$, and determine its quadratic variation. The continuity follows from $\Phi(C(\mathbb{R}_{+})_{0})=\{f\in C_{\kappa}(\mathbb{R}_{+}):f-g\in C(\mathbb{R}_{+})_{0}\}$, [25, Corollary 3.6.2 (iii)] and ¯$\overline{\mathbb{P}}(\Phi(C(\mathbb{R}_{+})_{0}))=\overline{\mathbb{P}}_{g}(C(\mathbb{R}_{+})_{0})=1.$ Let $s<t$, $m\in\mathbb{N}$, $f\colon\mathbb{R}^{m}\longrightarrow\mathbb{R}$ bounded and continuous, and $0\leq s_{1}\leq\dots\leq s_{m}\leq s$. Let $(x(t))_{t\geq 0}$ denote the coordinate process $x(t):C_{\kappa}(\mathbb{R}_{+};\mathbb{R})\longrightarrow\mathbb{R},\ \ w\longmapsto w(t)$. Applying [11, Theorem 3.5] yields
(14) $\mathbb{E}^{\overline{\mathbb{P}}}\left[f(x(s_{1}),\dots,x(s_{m}))\left(\mathcal{M}_{t}^{\Phi}-\mathcal{M}_{s}^{\Phi}\right)\right]=\lim_{n\longrightarrow\infty}\mathbb{E}[f(X^{n}_{s_{1}},\dots,X^{n}_{s_{m}})(M_{t}^{n}-M_{s}^{n})]=0.$
@@
where $\lambda=q-2>0$ if $g\in L^{q}_{\mathrm{loc}}(\mathbb{R}_{+}),q\in(2,\infty),$ and $\lambda=1$ if $g\in L^{\infty}_{\mathrm{loc}}(\mathbb{R}_{+})$. Thus, letting first $n\longrightarrow\infty$ and then $R\longrightarrow\infty$, proves the asserted. Arguing as in (14) yields that $(\mathcal{M}^{\Phi})^{2}-\int_{0}^{\cdot}\sigma(s,x(s))^{2}\,\mathrm{d}s$ is an $\overline{\mathbb{P}}$-martingale. Consequently, we have shown that $(\mathcal{M}^{\Phi}_{t})_{t\geq 0}$ is a continuous $\overline{\mathbb{P}}$-martingale with quadratic variation
@@
By martingale representation theorem [34, Chapter V, Theorem 3.9], there exists, possibly on an enlargement of the probability space, a Brownian motion $W$ and a predictable process $\alpha$ such that $\alpha_{s}^{2}=\sigma(s,\widehat{X}_{s})^{2}$ and $\mathcal{M}^{\Phi}_{t}=\int_{0}^{t}\alpha_{s}\,\mathrm{d}W_{s}$ where $\widehat{X}$ denotes the processes $x$ considered on the enlargement such that $\mathcal{L}(\widehat{X})=\mathcal{L}(x)$. Consequently, by [4, Lemma 2.6]
@@
Remark that $g$ given as in (9) automatically satisfies $g\in L^{q}_{\mathrm{loc}}(\mathbb{R}_{+})$ and hence condition (4) is satisfied whenever
@@
Next, we turn to the comparison principle beyond Lipschitz continuous coefficients. Since a comparison principle as formulated in Theorem 2.4 would imply pathwise uniqueness of (1), it is in this generality out of reach. However, by approximation, we show that for any two weak solutions $X^{1},X^{2}$ there exists a coupling $(\widehat{X}^{1},\widehat{X}^{2})$ on another filtered probability space on which the comparison principle holds.
@@
Let $(g_{i},K,b_{i},\sigma)$, $i=1,2,$ satisfy Assumption 2.2 with $b_{1}(t,x)\leq b_{2}(t,x)$, $K\colon(0,\infty)\longrightarrow\mathbb{R}_{+}$ is continuous, preserves non-negativity and suppose there exists $\delta\in(0,1/2)$ with (9) holds for $g_{1},g_{2}$. Moreover, assume that $g_{1},g_{2}$ satisfy (5) and are admissible. Then for any two continuous weak solutions $(X^{i},(\Omega^{i},\mathcal{F}^{i},\mathbb{F}^{i},\mathbb{P}^{i}),B^{i})$, $i=1,2,$ of (10), there exists a continuous weak solution $((\widehat{X}^{1},\widehat{X}^{2}),(\widehat{\Omega},\widehat{\mathcal{F}},\widehat{\mathbb{F}},\widehat{\mathbb{P}}),\widehat{B})$ of the two dimensional stochastic Volterra equation
@@
such that $\mathcal{L}(\widehat{X}^{i})=\mathcal{L}(X^{i})$, $i=1,2,$ and ^$\widehat{\mathbb{P}}[\widehat{X}^{1}_{t}\leq\widehat{X}^{2}_{t},\,t>0]=1$.
@@
Let us define $b_{1,n},b_{2,n},\sigma_{n}$ as in (11). Since $b_{1}\leq b_{2}$, it is clear that this approximation inherits the ordering condition $b_{1,n}\leq b_{2,n}$. Let $(X^{n})_{n\in\mathbb{N}},(Y^{n})_{n\in\mathbb{N}}$ be the family of processes constructed for $(g_{i},K,b_{i,n},\sigma_{n})$ on some fixed filtered probability space $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$. By Theorem 2.4, $\mathbb{P}[X^{n}_{t}\leq Y^{n}_{t},t>0]=1$. Following the arguments presented in the proof of Theorem 2.6, we firstly deduce that $((X^{n}-g_{1},Y^{n}-g_{2}))_{n\in\mathbb{N}}$ is a tight sequence of processes and hence has a weakly convergent subsequence on the path space $C(\mathbb{R}_{+};\mathbb{R}^{2})$. Secondly, we extend the spaces $C(\mathbb{R}_{+})_{0}\coloneqq C(\mathbb{R}_{+};\mathbb{R})_{0}$, $C_{\kappa}(\mathbb{R}_{+})\coloneqq C_{\kappa}(\mathbb{R}_{+};\mathbb{R})$ in an obvious way to $C(\mathbb{R}_{+};\mathbb{R}^{2})_{0}$, $C_{\kappa}(\mathbb{R}_{+};\mathbb{R}^{2})$ with $\kappa>\delta$ and infer that $((X^{n},Y^{n}))_{n\in\mathbb{N}}$ is weakly convergent on $C_{\kappa}(\mathbb{R}_{+};\mathbb{R}^{2})$ with weak limit denoted by $\overline{\mathbb{P}}.$ To prove the desired comparison result it is essential to observe that $\overline{\mathbb{P}}$ has inherited the ordering structure. Indeed, let $(x,y)=((x,y)(t))_{t\geq 0}$ be the coordinate process

### v2

@@ 2. Non-negativity preserving case
admits a unique, strong continuous solution for $i=1,2$. Since completely monotone kernels are non-negativity preserving and satisfy (3), see [7, Theorem 2.3], we may apply steps 1 and 2 to obtain $\mathbb{P}[X^{1,H}_{T}\leq X^{2,H}_{T}]=1$ for each $T>0$. Next, we prove that
@@
where $R$ denotes the resolvent of the second kind of $c(K^{H})^{2}$, and $c$ is some large enough constant. Finally, the first term can be bounded by
$\int_{0}^{T}|K(s)-K^{H}(s)|^{2+\eta}\,\mathrm{d}s$ $\leq\int_{0}^{T}\left(\int_{[H,\infty)}\mathrm{e}^{-\rho s}\,\mu(\mathrm{d}\rho)\right)^{2+\eta}\,\mathrm{d}s$
@@
Note that the case where $g_{i}(t)=x_{i}$ for $i=1,2$ was already covered in [7]. Since the fractional Riemann-Liouville kernel satisfies Assumption 2.1 and is completely monotone, Theorem 2.5 can be applied to this case.
@@
Let $X^{1},X^{2}$ be fractional Ornstein-Uhlenbeck processes given as the unique strong solutions of
@@
Since $E_{\alpha,\alpha}(\beta t^{\alpha})\geq 0$ for any $\beta,t$, the comparison principle is satisfied if and only if $E_{\alpha,\gamma}(\beta t^{\alpha})\geq 0$ for all $t\geq 0$. If $\beta\geq 0$, then $E_{\alpha,\gamma}(\beta t^{\alpha})\geq 0$ and nothing needs to be shown. Suppose that $\beta<0$. If $\gamma\geq\alpha$, then $t\longmapsto E_{\alpha,\gamma}(-t^{\alpha})$ is completely monotone, see [20, Section 4.9.2], and hence the comparison principle holds. Conversely, suppose that $\gamma<\alpha$. Then $(0,\infty)\ni t\longmapsto t^{\gamma-1}E_{\alpha,\gamma}(-|\beta|t^{\alpha})$ is integrable and its Laplace transform is given by $\int_{0}^{\infty}t^{\gamma-1}E_{\alpha,\gamma}(-|\beta|t^{\alpha})\mathrm{e}^{-st}\,\mathrm{d}s=\frac{s^{\alpha-\gamma}}{s^{\alpha}+|\beta|}$ for $s\geq 0$. Evaluating this in $s=0$ gives
$\int_{0}^{\infty}t^{\gamma-1}E_{\alpha,\gamma}(-|\beta|t^{\alpha})\,\mathrm{d}s=0$
and hence $(0,\infty)\ni t\longmapsto E_{\alpha,\gamma}(-|\beta|t^{\alpha})$ needs to attain negative values. Consequently, the comparison principle fails to hold. ∎
While the above proof is direct, one may also seek to apply Theorem 2.5 instead. Indeed, consider the fractional Riemann-Liouville kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha\in(1/2,1)$ and define $g_{i}(t)=x_{i}\frac{t^{\gamma-1}}{\Gamma(\gamma)}$ where we additionally assume that $\gamma\geq\alpha$. Then $\widetilde{g}_{i}(t)=0$ and $h_{i}(t)=x_{i}\frac{t^{(\gamma-\alpha)-1}}{\Gamma(\gamma-\alpha)}$. In particular, condition (ii) from Definition 2.4 reduces to
$(b_{2}-b_{1})+(x_{2}-x_{1})\left(\beta+\frac{t^{(\gamma-\alpha)-1}}{\Gamma(\gamma-\alpha)}\right)\geq 0.$
Hence, Theorem 2.5 is applicable for $\beta\geq 0$, while for $\beta<0$ it is generally not applicable.
Next, let us examine cases where the coefficients $b_{1},b_{2},\sigma$ are not Lipschitz continuous. Since, in this generality, the weak existence of solutions is not guaranteed from the literature, below we complement the latter by proving the weak existence of continuous solutions when the function $g$ may be singular in $t=0$. Our proof below employs some tightness arguments. However, in contrast to existing arguments such as [4, 26], here we need to introduce an auxiliary weighted space of continuous functions to control the order of singularity of the function $g$.
@@
then the stochastic Volterra equation (1) admits a continuous weak solution.
@@
Following [26, Lemma 3.6], we define for $f\in\{b,\sigma\}$ a sequence of functions
(9) $f_{n}(t,x)=\psi_{n}(x)\int_{\mathbb{R}}f(t,x-y)\varphi_{n}(y)\,\mathrm{d}y,\quad n\in\mathbb{N},$
where $\psi_{n}\colon\mathbb{R}\longrightarrow[0,1]$ with $\operatorname{supp}\psi_{n}\subset[-(n+1),n+1]$ and $\psi_{n}(x)=1$ for $x\in[-n,n]$, where $\varphi(y)=\frac{1}{c_{n}}(1-y^{2})^{n}\mathbbm{1}_{[-1,1]}(y)$ with $c_{n}=\int_{[-1,1]}(1-y^{2})^{n}\,\mathrm{d}y$. Then $(f_{n})_{n\in\mathbb{N}}$ is a sequence of Lipschitz continuous functions in the second variable and $\lim_{n\to\infty}f_{n}(t,\cdot)=f(t,\cdot)$ uniformly on compacts for all $t\geq 0$. Moreover, the following growth condition is satisfied uniformly in $n$,
(10) $|b_{n}(t,x)|+|\sigma_{n}(t,x)|\leq C(1+|x|)^{\xi},\quad\text{a.a. }t\geq 0,x\in\mathbb{R}.$
@@
holds for some $\theta>0$, $p\geq 2$ and a constant $c$ that only depends on $C$, $T$, $K$, pp. By applying Kolmogorov’s tightness criterion (see e.g. [29, Chapter XIII, Theorem 1.8]), we conclude that $(X^{n}-g)_{n\in\mathbb{N}}$ is a tight sequence of processes and hence has a weakly convergent subsequence on the path space $C(\mathbb{R}_{+})$ with limit denoted by $\overline{\mathbb{P}}_{g}$. By abuse of notation, we denote this subsequence also by $(X^{n}-g)_{n\in\mathbb{N}}$. Let $C(\mathbb{R}_{+})_{0}\subset C(\mathbb{R}_{+})$ be the space of continuous functions that vanish at $t=0$. Since $\mathcal{L}(X^{n}-g)(C(\mathbb{R}_{+})_{0})=1$ it follows by the Portmanteau theorem that
@@
Next, to prove weak convergence of $(X^{n})_{n\in\mathbb{N}}$, let $C_{\kappa}([0,T])$ with $\kappa>0$ be the Banach space of continuous functions $f\colon(0,T]\longrightarrow\mathbb{R}$ with finite norm $\|f\|_{\kappa,T}\coloneqq\sup_{t\in(0,T]}t^{\kappa}|f(t)|$ such that $\lim_{t\downarrow 0}t^{\kappa}f(t)$ exists. Note that $\iota\colon C_{\kappa}([0,T])\longrightarrow C([0,T])$ defined by $\iota(f)(t)=t^{\kappa}f(t)$ satisfies $\|\iota(f)\|_{[0,T]}=\|f\|_{\kappa,T}$, and hence is an isometric isomorphism, whence $C_{\kappa}([0,T])$ is a polish space. Finally, we let $C_{\kappa}(\mathbb{R}_{+})\coloneqq\bigcap_{k\in\mathbb{N}}C_{\kappa}([0,k])$ be equipped with the metric
@@
By assumption (8), $g\in C_{\kappa}(\mathbb{R}_{+})$ for $\kappa>\delta$. Hence $\Phi\colon C(\mathbb{R}_{+})_{0}\longrightarrow C_{\kappa}(\mathbb{R}_{+})$ given by $\Phi(f)=f+g$ defines a continuous mapping. Thus, $(\Phi(X^{n}-g))_{n\in\mathbb{N}}=(X^{n})_{n\in\mathbb{N}}$ is weakly convergent on $C_{\kappa}(\mathbb{R}_{+})$ with weak limit denoted by $\overline{\mathbb{P}}$. Note that $\overline{\mathbb{P}}_{g},\overline{\mathbb{P}}$ are related by
(11) $\overline{\mathbb{P}}_{g}\circ\Phi^{-1}=\overline{\mathbb{P}}.$
It remains to show that $\overline{\mathbb{P}}$ determines a weak solution of (1).
Let $L$ be the resolvent of the first kind of KK111i.e. a locally finite (positive) measure $L$ such that $L\ast K=K\ast L=1$. The latter exists since $K\neq 0$ is non-negative and non-increasing, see [21, Theorem 5.5.5].. Let $(M_{t}^{n})_{t\geq 0}$ be the continuous $\mathbb{P}$-martingale for the filtration $\mathbb{F}$ given by
@@
where the continuity is guaranteed by [21, Corollary 3.6.2 (iii)]. Since $X^{n}-g\in C(\mathbb{R}_{+})_{0}$, by continuous mapping theorem combined with (11), we obtain
@@
Next, we show that $(\mathcal{M}^{\Phi}_{t})_{t\geq 0}$ is a continuous $\overline{\mathbb{P}}$-martingale for the natural filtration generated by coordinate process on $C_{\kappa}(\mathbb{R}_{+})$, and determine its quadratic variation. The continuity follows from $\Phi(C(\mathbb{R}_{+})_{0})=\{f\in C_{\kappa}(\mathbb{R}_{+}):f-g\in C(\mathbb{R}_{+})_{0}\}$, [21, Corollary 3.6.2 (iii)] and $\overline{\mathbb{P}}(\Phi(C(\mathbb{R}_{+})_{0}))=\overline{\mathbb{P}}_{g}(C(\mathbb{R}_{+})_{0})=1.$ Let $s<t$, $m\in\mathbb{N}$, $f\colon\mathbb{R}^{m}\longrightarrow\mathbb{R}$ bounded and continuous, and $0\leq s_{1}\leq\dots\leq s_{m}\leq s$. Let $(x(t))_{t\geq 0}$ denote the coordinate process $x(t)\colon C_{\kappa}(\mathbb{R}_{+};\mathbb{R})\longrightarrow\mathbb{R},\ \ w\longmapsto w(t)$. Applying [11, Theorem 3.5] yields
(12) $\mathbb{E}^{\overline{\mathbb{P}}}\left[f(x(s_{1}),\dots,x(s_{m}))\left(\mathcal{M}_{t}^{\Phi}-\mathcal{M}_{s}^{\Phi}\right)\right]=\lim_{n\to\infty}\mathbb{E}[f(X^{n}_{s_{1}},\dots,X^{n}_{s_{m}})(M_{t}^{n}-M_{s}^{n})]=0.$
@@
where $\lambda=q-2>0$ if $g\in L^{q}_{\mathrm{loc}}(\mathbb{R}_{+}),q\in(2,\infty),$ and $\lambda=1$ if $g\in L^{\infty}_{\mathrm{loc}}(\mathbb{R}_{+})$. Thus, letting first $n\longrightarrow\infty$ and then $R\longrightarrow\infty$, proves the assertion. Arguing as in (12) yields that $(\mathcal{M}^{\Phi})^{2}-\int_{0}^{\cdot}\sigma(s,x(s))^{2}\,\mathrm{d}s$ is an $\overline{\mathbb{P}}$-martingale. Consequently, we have shown that $(\mathcal{M}^{\Phi}_{t})_{t\geq 0}$ is a continuous $\overline{\mathbb{P}}$-martingale with quadratic variation
@@
By martingale representation theorem [29, Chapter V, Theorem 3.9], there exists, possibly on an enlargement of the probability space, a Brownian motion $W$ and a predictable process $\alpha$ such that $\alpha_{s}^{2}=\sigma(s,\widehat{X}_{s})^{2}$ and $\mathcal{M}^{\Phi}_{t}=\int_{0}^{t}\alpha_{s}\,\mathrm{d}W_{s}$ where $\widehat{X}$ denotes the processes $x$ considered on the enlargement such that $\mathcal{L}(\widehat{X})=\mathcal{L}(x)$. Consequently, by [4, Lemma 2.6]
@@
Remark that $g$ given as in (8) automatically satisfies $g\in L^{q}_{\mathrm{loc}}(\mathbb{R}_{+})$, $q\in(0,1/\delta)$, and hence condition (2) is satisfied whenever
@@
As another consequence, we may also prove the existence of nonnegative continuous weak solutions. The latter complements the results obtained in [2].
Corollary 2.8.
Let $K\colon(0,\infty)\longrightarrow\mathbb{R}_{+}$ be continuous and preserve nonnegativity, and let $(g,K,b,\sigma)$ satisfy Assumption 2.1. Moreover, suppose that there exists a sequence of globally Lipschitz continuous functions $b_{n},\sigma_{n}\colon\mathbb{R}_{+}\times\mathbb{R}\longrightarrow\mathbb{R}$ with property (10) such that $b_{n}(t,\cdot)\longrightarrow b(t,\cdot)$, $\sigma_{n}(t,\cdot)\longrightarrow\sigma(t,\cdot)$ locally uniformly for each $t\geq 0$, and it holds that
(13) $b_{n}(t,0)\geq 0\ \text{ and }\ \sigma_{n}(t,0)=0,\qquad t\geq 0,\ n\in\mathbb{N}.$
If $g=\widetilde{g}+K\ast h$ is such that $0\leq\widetilde{g}\in C((0,\infty))\cap L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})$ is nondecreasing, $0\leq h\in L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})$, (5),(8) hold for some $\delta\in(0,1/2)$, then the stochastic Volterra equation (1) admits a nonnegative continuous weak solution.
Proof.
Let $X^{n}$ be the unique strong solution of (1) with $b,\sigma$ replaced by $b_{n},\sigma_{n}$, i.e.
$X_{t}^{n}=g(t)+\int_{0}^{t}K(t-s)b_{n}(s,X_{s}^{n})\,\mathrm{d}s+\int_{0}^{t}K(t-s)\sigma_{n}(s,X_{s}^{n})\,\mathrm{d}B_{s}.$
By the convergence arguments given in the proof of Theorem 2.7, it suffices to prove that $\mathbb{P}[X^{n}_{t}\geq 0,\,t>0]=1$. To prove the latter, we apply Theorem 2.5 to the pair of processes $X^{n}$ and $Y\equiv 0$. Indeed, note that $Y\equiv 0$ is the unique solution of
$Y_{t}^{n}=\widehat{g}_{n}(t)+\int_{0}^{t}K(t-s)b_{n}(s,Y^{n}_{s})\,\mathrm{d}s+\int_{0}^{t}K(t-s)\sigma_{n}(s,Y^{n}_{s})\,\mathrm{d}B_{s}$
where $\widehat{g}_{n}(t)=-\int_{0}^{t}K(t-s)b_{n}(s,0)\,\mathrm{d}s$. It is easy to check that $(g,K,b_{n},\sigma_{n})$ and $(\widehat{g}_{n},K,b_{n},\sigma_{n})$ are comparable in the sense of Definition 2.4. Hence Theorem 2.5 gives $\mathbb{P}[X^{n}\geq Y=0,\,t>0]=1$, which proves the assertion. ∎
Below, we illustrate this existence result by an application to a general one-dimensional nonnegative process with a power-law diffusion coefficient. The latter contains the rough Cox-Ingersoll-Ross process as a particular case.
Example 2.9.
Let $\lambda,\theta,\sigma>0$, $\gamma_{1},\gamma_{2}\in(0,1]$, and $\alpha\in(1/2,1)$. Take $g$ of the form
$g(t)=\widetilde{g}(t)+\int_{0}^{t}\frac{(t-s)^{\alpha-1}}{\Gamma(\alpha)}h(s)\,\mathrm{d}s$
such that $0\leq\widetilde{g}\in C((0,\infty))\cap L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})$ is nondecreasing, $0\leq h\in L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})$, and (5), (8) hold for some $\delta\in(0,1/2)$. If Assumption 2.1 is satisfied for
$X_{t}=g(t)+\int_{0}^{t}\frac{(t-s)^{\alpha-1}}{\Gamma(\alpha)}\lambda\left(\theta-X_{s}\right)^{\gamma_{1}}\,\mathrm{d}s+\sigma\int_{0}^{t}\frac{(t-s)^{\alpha-1}}{\Gamma(\alpha)},X_{s}^{\gamma_{2}}\,\mathrm{d}B_{s},$
then it admits a continuous nonnegative weak solution.
Proof.
The assertion follows from Corollary 2.8, once we have verified condition (13). Here, we may take $\sigma_{n}(x)=n^{1-\gamma}|x|\mathbbm{1}_{\{|x|\leq\frac{1}{n}\}}+|x|^{\gamma}\mathbbm{1}_{\{|x|>\frac{1}{n}\}}$ for $\sigma(x)=|x|^{\gamma}$ and similarly for the drift coefficient. Then it is easy to see that (13) holds, which proves the assertion. ∎
Next, we turn to the comparison principle beyond Lipschitz continuous coefficients. Since a comparison principle as formulated in Theorem 2.5 would imply pathwise uniqueness of (1), it is in this generality out of reach. However, by approximation, we show that there always exists a couple of solutions $(X^{1},X^{2})$ defined on a joint filtered probability space that satisfies the comparison principle, i.e., the existence of a monotone coupling.
@@
[…]
@@
such that $\widehat{\mathbb{P}}[\widehat{X}^{1}_{t}\leq\widehat{X}^{2}_{t},\,t>0]=1$. In particular, its marginals are continuous weak solutions of (1).
@@
Let us define $b_{1,n},b_{2,n},\sigma_{n}$ as in (9). Since $b_{1}\leq b_{2}$, it is clear that this approximation inherits the ordering condition $b_{1,n}\leq b_{2,n}$. Let $(X^{n})_{n\in\mathbb{N}},(Y^{n})_{n\in\mathbb{N}}$ be the family of processes constructed for $(g_{i},K,b_{i,n},\sigma_{n})$ on some fixed filtered probability space $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$. By Theorem 2.5, $\mathbb{P}[X^{n}_{t}\leq Y^{n}_{t},t>0]=1$.
Following the arguments presented in the proof of Theorem 2.7, we firstly deduce that $((X^{n}-g_{1},Y^{n}-g_{2}))_{n\in\mathbb{N}}$ is a tight sequence of processes and hence has a weakly convergent subsequence on the path space $C(\mathbb{R}_{+};\mathbb{R}^{2})$. Secondly, we extend the spaces $C(\mathbb{R}_{+})_{0}\coloneqq C(\mathbb{R}_{+};\mathbb{R})_{0}$, $C_{\kappa}(\mathbb{R}_{+})\coloneqq C_{\kappa}(\mathbb{R}_{+};\mathbb{R})$ in an obvious way to $C(\mathbb{R}_{+};\mathbb{R}^{2})_{0}$, $C_{\kappa}(\mathbb{R}_{+};\mathbb{R}^{2})$ with $\kappa>\delta$ and infer that $((X^{n},Y^{n}))_{n\in\mathbb{N}}$ is weakly convergent on $C_{\kappa}(\mathbb{R}_{+};\mathbb{R}^{2})$ with weak limit denoted by $\overline{\mathbb{P}}.$ To prove the desired comparison result, it is essential to observe that $\overline{\mathbb{P}}$ has inherited the ordering structure. Indeed, let $(x,y)=((x,y)(t))_{t\geq 0}$ be the coordinate process